OpenStudy (anonymous):

Help plz will award medal and become fan The bathtub in Tanya's bathroom is clogged up, and she needs to empty out the water in it. The only thing she can find to use is a bucket that holds of a gallon of water. If it takes 30 full buckets to empty the tub, how much water was in the tub originally? 25 gallons 33 gallons 27 gallons 32 gallons

OpenStudy (anonymous):

The bucket can contain 9/10 of a gallon of water.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\text{number of buckets}}{\text{amount of water}}=\frac{\text{number of buckets}}{\text{amount of water}}\] \[\Large \frac{1 \ \text{bucket}}{9/10 \ \text{gallons}}=\frac{30 \ \text{buckets}}{x \ \text{gallons}}\] \[\Large \frac{1}{9/10}=\frac{30}{x}\] \[\Large 1*x=(9/10)*30\] \[\Large x=??\] I'll let you finish up

OpenStudy (anonymous):

Is it 27 buckets.

jimthompson5910 (jim_thompson5910):

yeah because the `(9/10)*30` turns into `27`

OpenStudy (anonymous):

ok Thank you.